OpenStudy (anonymous):

Multiply and simplify y/16 *4y^2/y^2

OpenStudy (anonymous):

I think your expression is this: \[\frac{ y }{ 16 } \times \frac{4y^2}{y^2}\] If you noticed, those y^2 are cancelled out: \[\frac{y}{16} \times 4 = \frac{y}{4}\]